Option Informatique en Spé MP^{∗} — Année scolaire 2007/2008

1 BESSON Étienne
2 BIESSY Guillaume
3 BOISSON Damien
4 CAVAILLOLES David*
5 CHAIGNEAU Matthieu
6 CORLAY Guillaume
7 CULAN Christophe
8 DE MARIGNAN Louis
9 GUELFUCCI Jérôme
10 HILICO Adèle
11 HOONAKER Raphaël
12 LANUSSE François
13 LETURC Kevin
14 PUCHOL Martin
15 RAYNEL Kevin
16 RENAUDIN Rafael
17 THOU Sutkhemaroath*